OpenStudy (anonymous):

What is the local maximum value of this function? (Round answer to the nearest thousandth.) g(x)=x^3+5x^2-17x-21

OpenStudy (anonymous):

You need to find the critical points, which means the points at which the derivative = 0. So find the derivative: g(x)=x^3+5x^2-17x-21 g'(x) = 3x^2 + 10x - 17 Set equal to 0: 3x^2 + 10x - 17 Find a, b such that ab = -17, a + b = 10 and factor. Post your work.